# A Diffusion-based Xray2MRI Model: Generating Pseudo-MRI Volumes From one Single X-ray

### v2

@@ Abstract
Knee osteoarthritis (KOA) is a prevalent musculoskeletal disorder, and X-rays are commonly used for its diagnosis due to their cost-effectiveness. Magnetic Resonance Imaging (MRI), on the other hand, offers detailed soft tissue visualization and has become a valuable supplementary diagnostic tool for KOA. Unfortunately, the high cost and limited accessibility of MRI hinders its widespread use, leaving many patients with KOA to rely solely on X-ray imaging. In this study, we introduce a novel diffusion-based Xray2MRI model capable of generating pseudo-MRI volumes from a single X-ray image. In addition to using X-rays as conditional input, our model integrates target depth, KOA probability distribution, and image intensity distribution modules to guide the synthesis process, ensuring that the generated corresponding slices accurately correspond to the anatomical structures. Experimental results demonstrate that by integrating information from X-rays with additional input data, our proposed approach is capable of generating pseudo-MRI sequences that approximate real MRI scans. In addition, by increasing the number of inference steps, the model achieves effective interpolation, which further improves the continuity and smoothness of the generated MRI sequences, representing a promising first attempt at cost-effective medical imaging solutions. This study is available on https://zwang78.github.io/.
Keywords Knee osteoarthritis  $\cdot$ X-ray  $\cdot$ Magnetic Resonance Imaging  $\cdot$ Diffusion-based
1 Introduction
@@
To address these limitations, researchers have explored various techniques to enhance MRI reconstruction and reduce its dependency on traditional methods. Fessler et al. [12] provides an in-depth review of iterative algorithms used for improving MRI image reconstruction, particularly in scenarios where traditional inverse Fourier transforms are inadequate. The author highlights the limitations of conventional methods, especially when dealing with non-Cartesian sampling, under-sampled data, or the presence of nonlinear magnetic fields, and then focuses on model-based approaches, which utilize iterative methods that incorporate physical models of MRI acquisition, allowing for better image quality by addressing inhomogeneity and under-sampling issues. The authors demonstrate that model-based methods can significantly enhance image reconstruction, providing superior results compared to traditional techniques, though often at the cost of increased computation. In [13], Han et al. proposed leveraging the k-space to improve the efficiency and accuracy of MRI reconstruction. The authors introduced a framework inspired by the low-rank Hankel matrix completion approach, which directly interpolates missing k-space data and enhances the representation of MRI data by learning the underlying signal structure, resulting in high-quality images even from undersampled data. The proposed k-space deep learning approach significantly reduces the computational complexity and memory requirements. In [14], Lustig et al. leveraged compressed sensing to significantly accelerate MRI scan times without compromising image quality. The authors explore the inherent sparsity of MR images, both in the image domain and in transform domains. By randomly undersampling k-space and using nonlinear recovery schemes, the study shows that high-quality MR images can be reconstructed from incomplete data. This approach effectively reduces acquisition time and mitigates artefacts that arise from undersampling, which demonstrates that compressed sensing has the potential to substantially reduce scan times in clinical settings, making MRI more efficient and accessible.
@@
Inspired by the above work, we propose a novel approach named Xray2MRI, based on the Conditional Latent Diffusion Model (CLDM) [20]. Specifically, during the training phase, each real MRI slice is first noised and then progressively denoised using the corresponding X-ray as conditional input and the adaptive-weight guidance of the target depth, image intensity distribution, and KOA probability distribution to reconstruct the input MRI slice. In the inference phase, given a knee X-ray and a target depth, the model progressively denoises from Gaussian noise to generate the corresponding MRI slice at that depth. Then, the target depth is adjusted and the denoising process is repeated iteratively for each depth until all the required slices are synthesized. These individual slices are then stacked together to form a full 3D MRI volume. It is noteworthy that the reconstructed MRI is not intended to serve as a clinical replacement for a real MRI. Still, it marks a pioneering cross-modal attempt to generate corresponding pseudo-MRI volumes from X-rays, for seeking cost-effective medical imaging solutions.
The primary contributions of this paper include the following:
@@
    A novel CLDM-based framework, named Xray2MRI, is proposed to generate pseudo-MRI volumes from a single X-ray image.
@@
    In addition to the target depth, information about KOA probability distribution and image intensity distribution from the X-ray is combined via an adaptive-weight module to guide the denoising process.
@@
    A new SSIM-based evaluation metric is specially developed. The multi-indicator assessment approach ensures both quantitative and qualitative validation of the synthesized MRI images for both image quality and clinical relevance.
@@
    The model demonstrates effective interpolation between MRI slices by increasing the number of inference steps (i.e., exceeding the depth of the original MRI volume), resulting in smoother transitions and improved continuity between the generated slices.
2 Proposed approach
2.1 Classical conditional latent diffusion model
Before delving into the proposed approach, we provide a brief overview of the classic conditional latent diffusion model, which serves as the foundation for our methodology.
@@ 2.1 Classical conditional latent diffusion model
Figure 2: The flowchart of the proposed approach, starts by extracting a slice at position $d$ from a real MRI sequence $x$, denoted as $x_{d}$. This slice is encoded by the encoder $\mathcal{E}_{1}$ to produce a latent representation $z_{0}$, which is progressively noised following the distribution $q(z_{t}|z_{t-1})$, ultimately generating the noised latent representation $z_{T}$. The corresponding X-ray image $x_{c}$ is then used as a conditional input, encoded by $\mathcal{E}_{2}$, and concatenated with $z_{T}$ to form the initial input for the downsampling network. During the denoising process, the target depth $d$ is embedded at various stages of the downsampling phase. Concurrently, the KOA probability distribution $p$, obtained from a pre-trained KOA classifier $\mathcal{C}$, and the image intensity distribution $i$ from the X-ray image are integrated into the upsampling phase through an adaptive weighting module, providing dynamic guidance. This denoising process is repeated for $T-1$ iterations to progressively refine the latent representation, resulting in the denoised latent $\hat{z}_{0}$, which is decoded by $\mathcal{D}$ to reconstruct the MRI slice $\hat{x}_{d}$ at depth $d$. During inference, the process begins with $z_{T}$ and is repeated $S$ times, corresponding to the number of channels $S$ in the original MRI sequence. Finally, the full pseudo-MRI sequence $\hat{x}$ is obtained by stacking all the generated MRI slices in sequential order (i.e., $\hat{x}$ = $\{\hat{x}_{d=1}$, $\hat{x}_{d=2}$,…,$\hat{x}_{d=S}\}$).
As shown in Fig. 1, a conditional latent diffusion model is a type of generative model used for producing data (such as images, audio, or text) based on certain conditions or inputs. This model combines the strengths of diffusion models and latent variable models, allowing it to generate complex data distributions while being guided by specific conditions. Different modules are to be introduced, respectively.
2.1.1 Encode process
@@ 2.1.2 Diffusion process
Here, we directly sample $z_{t}$ for any step $t$ based on the original data $z_{0}$: $z_{t}\sim q(z_{t}\mid z_{0})$ defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$. Through the reparameterization technique, we have:
@@ 2.1.3 Denoising process
$p_{\theta}(z_{t-1}\mid z_{t},x_{c})=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,x_{c}),\sigma_{\theta}^{2}(t))$ (5)
where $p_{\theta}(z_{t-1}\mid z_{t},x_{c})$ represents parameterized Gaussian distributions, the mean $\mu_{\theta}(z_{t},t,x_{c})$ and variance $\sigma_{\theta}^{2}(t)$ at each time step with the condition $x_{c}$ are determined by trained learning networks.
2.1.4 Training process
@@
$\mathcal{L}_{diff}=\mathbb{E}_{z_{0}\sim\mathcal{E}(x),x_{c},t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(z_{t},t,x_{c})||_{2}^{2}$ (7)
where $\mathbb{E}$ represents the expectation over all data points under the joint distribution of the conditioning information $x_{c}$, time steps $t$, and Gaussian noise $\epsilon$.
2.2 Proposed approach
@@
Conditional image $x_{c}$ is a corresponding X-ray image, which is then encoded by the encoder $\mathcal{E}_{2}$ and concatenated with the $z_{T}$ as the input of the U-Net, serves as the beginning of the reverse process. The employed 2D U-Net architecture adopts a multi-scale hierarchical structure comprising a time embedding module, an encoder module, a middle module, and a decoder module. This UNet model takes 8-channel input and produces 4-channel output, built upon a base channel configuration of 320. The encoder module begins with a 2D convolutional layer featuring a kernel size of 3 and padding of 1, followed by a series of residual blocks with progressively increasing feature dimensions. Each residual block includes multiple convolutional layers, GroupNorm [24] for normalization, the Swish activation function [25] for non-linearity, and optional dropout for regularization. Two residual blocks are employed per downsampling stage, with attention mechanisms activated at resolutions of 4, 2, and 1. The channel count increases across network layers following a multiplier sequence of 1, 2, 4, and 4. The decoder mirrors this structure, utilizing the same residual blocks to progressively upsample feature maps, reducing the number of channels in reverse order while merging them with corresponding encoder feature maps through skip connections. Additionally, the model incorporates 8 attention heads and employs a spatial transformer with a depth of 1, conditioned on a context dimension of 768.
Next, the target depth $d$ was embedded into the downsampling stage to allow the model to condition the latent representation with spatial hierarchy early in the process. This ensures that depth-related features, such as different tissue layers and bone structures, are incorporated from the outset, preserving the anatomical coherence of the volume. In the upsampling stage, the KOA probability distribution $p$ and image intensity distribution $i$ were embedded to guide the synthesis of fine clinical features based on the spatial hierarchy established during downsampling, such as the presence of osteophytes and the size of the joint space. However, since the variation between slices in MRI is not uniform, different slices have varying levels of sensitivity to the guided information. Some slices may require more emphasis on specific anatomical features, while others may need more focus on KOA severity. To address this, we introduce an automatic weight adjustment mechanism that dynamically combines the information from the KOA probability distribution $p$ and the image intensity distribution $i$ to ensure that the final MRI outputs maintain a higher level of accuracy and clinical relevance by properly leveraging the complementary strengths of both distributions. The adaptive-weight module $w$ begins by independently adjusting the weights of $p$ and $i$ using the $L^{2}$-norm, scaled by the learnable parameters $\mu_{p}$ and $\mu_{i}$, respectively:
$G_{p}=\mu_{p}\|p\|_{2}=\mu_{p}\left(\sum_{m=1}^{h}\sum_{n=1}^{w}(p_{m,n})^{2}+\kappa\right)^{\frac{1}{2}}$ (9)
$G_{i}=\mu_{i}\|i\|_{2}=\mu_{i}\left(\sum_{m=1}^{h}\sum_{n=1}^{w}(i_{m,n})^{2}+\kappa\right)^{\frac{1}{2}}$ (10)
where $h$ and $w$ represent the height and width of the feature maps $p$ and $i$. $p_{m,n}$ and $i_{m,n}$ represent the values of the respective feature maps at the pixel located at row $m$ and column $n$. $\kappa$ is a small constant added to prevent division by zero, typically set to an empirical value of 1e-05.
Next, the calculated weights $G_{p}$ and $G_{i}$ are normalized as follows:
$\hat{G}_{p}=\frac{G_{p}}{G_{p}+G_{i}+\epsilon},\quad\hat{G}_{i}=\frac{G_{i}}{G_{p}+G_{i}+\epsilon}$ (11)
where $\epsilon$ is a small positive constant (e.g., 1e-08) added for numerical stability, preventing division by zero.
Then, $p$ and $i$ are adjusted by applying the normalized weights:
$\tilde{p}=p\left[1+\sigma(\nu_{p}\hat{G}_{p}+o_{p})\right],\quad\tilde{i}=i\left[1+\sigma(\nu_{i}\hat{G}_{i}+o_{i})\right]$ (12)
where $\sigma$ is a Sigmoid function. $v_{p}$, $v_{i}$, and $o_{p}$, $o_{i}$ are learnable scaling and offset parameters to $p$ and $i$, respectively.
To capture the interaction between $p$ and $i$, a joint condition weight $\lambda_{pi}$ is learned using an attention mechanism [26], which is represented as:
$\lambda_{pi}=\text{Attention}(p,i)=\text{softmax}\left(\frac{(W_{p}p)^{\top}(W_{i}i)}{\sqrt{p_{k}}}\right)$ (13)
where $W_{p}$ and $W_{i}$ are learnable weight matrix associated with $p$ and $i$, respectively. $\sqrt{p_{k}}$ is a scaling factor.
The final combined condition $\tilde{y}_{pi}$ with the learned joint condition weight $\lambda_{pi}$ is as follows:
$\tilde{y}_{pi}=\lambda_{pi}\cdot(\tilde{p}+\tilde{i})$ (14)
$\mathop{\min}_{\theta}\mathbb{E}_{z\sim\varepsilon_{1}(x_{d}),t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}\{z_{t},t,[z_{T},\mathcal{C}(x_{c}),w(p,i)]\}||_{2}^{2}$ (15)
@@
Finally, for each reconstructed slice at the depth $d$, denoted as $\hat{x}_{d}$, it is generated through a decoder $\mathcal{D}$. During the inference phase, this process is repeated $S$ times, corresponding to the channel (depth) $S$ of the original MRI volume. The final generated pseudo-MRI sequence $\hat{x}$ is formed by stacking all generated slices, represented as $\hat{x}=\{\hat{x}_{0},\hat{x}_{1},\dots,\hat{x}_{s-1}\}$. For more clarity, Algorithm 1 and Algorithm 2 describe the training and inference process of the proposed approach, respectively.
Input: $x$, $t$, $S$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$ and all initial parameters
Output: $w$, $\hat{x}_{d}$, $\mathcal{E}_{1}$, $\mathcal{D}$, $\mathcal{U}$ and learned parameters
@@
             Get $p,i\leftarrow\mathcal{C}(x_{c}),intensity(x_{c})$
             Get $\epsilon_{\theta},\hat{z}_{d}\leftarrow\mathcal{U}(z_{T},\mathcal{E}_{2}(x_{c}),t,w(p,i))$
@@
Input: $t$, $w$, $S$, $x_{c}$, $\mathcal{E}_{2}$, $\mathcal{C}$, $\mathcal{D}$ and $\mathcal{U}$
@@
       Get $p,i\leftarrow\mathcal{C}(x_{c}),intensity(x_{c})$
       Get $\hat{z}_{d}\leftarrow\mathcal{U}(\epsilon,\mathcal{E}_{2}(x_{c}),t,w(p,i))$
@@
3 Experimental settings
3.1 Employed database

### v3

# Feasibility Study of a Diffusion-Based Model for Cross-Modal Generation of Knee MRI from X-ray: Integrating Radiographic Feature Information

@@ Abstract
Knee osteoarthritis (KOA) is a prevalent musculoskeletal disorder, often diagnosed using X-rays due to its cost-effectiveness. While Magnetic Resonance Imaging (MRI) provides superior soft tissue visualization and serves as a valuable supplementary diagnostic tool, its high cost and limited accessibility significantly restrict its widespread use. To explore the feasibility of bridging this imaging gap, we conducted a feasibility study leveraging a diffusion-based model that uses an X-ray image as conditional input, alongside target depth and additional patient-specific feature information, to generate corresponding MRI sequences. Our findings demonstrate that the MRI volumes generated by our approach is visually closer to real MRI scans. Moreover, increasing inference steps enhances the continuity and smoothness of the synthesized MRI sequences. Through ablation studies, we further validate that integrating supplementary patient-specific information, beyond what X-rays alone can provide, enhances the accuracy and clinical relevance of the generated MRI, which underscores the potential of leveraging external patient-specific information to improve the MRI generation. This study is available at https://zwang78.github.io/.
Index Terms:
Knee osteoarthritis, X-ray, Magnetic Resonance Imaging, Diffusion-based
I Introduction
@@
To address these limitations, researchers have explored various techniques to enhance MRI reconstruction and reduce its dependency on traditional methods. Fessler et al. [12] provides an in-depth review of iterative algorithms used for improving MRI image reconstruction, particularly in scenarios where traditional inverse Fourier transforms are inadequate. The author highlights the limitations of conventional methods, especially when dealing with non-Cartesian sampling, under-sampled data, or the presence of nonlinear magnetic fields, and then focuses on model-based approaches, which utilize iterative methods that incorporate physical models of MRI acquisition, allowing for better image quality by addressing inhomogeneity and under-sampling issues. The authors demonstrate that model-based methods can significantly enhance image reconstruction, providing superior results compared to traditional techniques, though often at the cost of increased computation. In [13], Han et al. proposed leveraging the k-space to improve the efficiency and accuracy of MRI reconstruction. The authors introduced a framework inspired by the low-rank Hankel matrix completion approach, which directly interpolates missing k-space data and enhances the representation of MRI data by learning the underlying signal structure, resulting in high-quality images even from undersampled data. The proposed k-space deep learning approach significantly reduces the computational complexity and memory requirements. In [14], Lustig et al. leveraged compressed sensing to significantly accelerate MRI scan times without compromising image quality. The authors explore the inherent sparsity of MR images, both in the image domain and in the transform domains. By randomly undersampling k-space and using nonlinear recovery schemes, the study shows that high-quality MR images can be reconstructed from incomplete data. This approach effectively reduces acquisition time and mitigates artefacts that arise from undersampling, which demonstrates that compressed sensing has the potential to substantially reduce scan times in clinical settings, making MRI more efficient and accessible.
@@
Inspired by the above work, we propose a novel approach based on the Conditional Latent Diffusion Model (CLDM) [20]. Specifically, during the training phase, each real MRI slice is first Gaussian noised and then progressively denoised using the corresponding X-ray as conditional input $y$, target depth $d$ and radiographic features $r$ of the specific patient as the guidance module to reconstruct the input MRI slice. In the inference phase, given a knee X-ray image, a target depth $d$ and the radiographic features $r$, the model progressively denoises from Gaussian noise to generate the corresponding MRI slice at that depth. Then, the target depth is adjusted and the denoising process is repeated iteratively for each depth until all the required slices are synthesized. These individual slices are then stacked together to form a fully generated 3D MRI volume. It is noteworthy that the reconstructed MRI is not intended to serve as a clinical replacement for a real MRI. Nevertheless, it signifies a groundbreaking feasibility study in integrating patient-specific radiographic information with knee X-rays to generate corresponding MRI volumes.
The primary contributions of this study are summarized as follows:
@@
    A novel CLDM-based framework is proposed for generating MRI volumes from X-ray images.
@@
    Beyond the target depth, patient-specific radiographic features not present in the X-ray are integrated as guidance to improve the denoising process.
@@
    Ablation studies demonstrate that incorporating additional patient-specific information enhances the accuracy and clinical relevance of the generated MRI.
@@
    By increasing the number of inference steps (i.e., exceeding the depth of the original MRI volume), the model achieves effective interpolation between MRI slices.
II Proposed approach
II-A Classical conditional latent diffusion model
Before delving into the proposed approach, we provide a brief overview of the classic conditional latent diffusion model, which serves as the foundation for our methodology. As shown in Fig. 1, a conditional latent diffusion model is a type of generative model used for producing data (such as images, audio, or text) based on certain conditions or inputs. This model combines the strengths of diffusion models and latent variable models, allowing it to generate complex data distributions while being guided by specific conditions. Different modules are to be introduced, respectively.
@@ II-A Classical conditional latent diffusion model
Figure 2: The flowchart of the proposed approach, starts by extracting a slice at position $d$ from a real MRI sequence $x$, denoted as $x_{d}$. This slice is encoded by the encoder $\mathcal{E}_{1}$ to produce a latent representation $z_{0}$, which is progressively noised following the distribution $q(z_{t}|z_{t-1})$, ultimately generating the noised latent representation $z_{T}$. The corresponding X-ray image $y$ is then used as a conditional input, encoded by $\mathcal{E}_{2}$, and concatenated with $z_{T}$ to form the initial input for the downsampling network. During the denoising process, the target depth $d$ and the patient-specific radiographic feature information $r$ are jointly embedded at various stages of the downsampling and upsampling phases, collectively providing denoising guidance. This denoising process is repeated for $T-1$ iterations to progressively refine the latent representation, resulting in the denoised latent $\hat{z}_{0}$, which is decoded by $\mathcal{D}$ to reconstruct the MRI slice $\hat{x}_{d}$ at depth $d$. During inference, the process begins with $z_{T}$ and is repeated $S$ times, corresponding to the number of channels $S$ in the original MRI sequence. Finally, the full generated MRI sequence $\hat{x}$ is obtained by stacking all the generated MRI slices in sequential order (i.e., $\hat{x}=\{\hat{x}_{d}\}_{d=0}^{S-1}$).
II-A1 Encode process
@@ II-A2 Diffusion process
Here, we directly sample $z_{t}$ for any step $t$ based on the original data $z_{0}$: $z_{t}\sim q(z_{t}\mid z_{0})$ defining $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}t=\prod{i=1}^{t}\alpha_{i}$. Through the reparameterization technique [22], we have:
@@ II-A3 Denoising process
$p_{\theta}(z_{t-1}\mid z_{t},y)=\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t,y),\sigma_{\theta}^{2}(t))$ (5)
where $p_{\theta}(z_{t-1}\mid z_{t},y)$ represents parameterized Gaussian distributions, the mean $\mu_{\theta}(z_{t},t,y)$ and variance $\sigma_{\theta}^{2}(t)$ at each time step with the condition $y$ are determined by trained learning networks.
II-A4 Training process
@@
$\mathcal{L}_{diff}=\mathbb{E}_{z_{0}\sim\mathcal{E}(x),y,t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}(z_{t},t,y)||_{2}^{2}$ (7)
where $\mathbb{E}$ represents the expectation over all data points under the joint distribution of the conditioning information $y$, time steps $t$, and Gaussian noise $\epsilon$.
II-B Proposed approach
@@
Conditional image $y$ is a corresponding knee joint X-ray image, which is then encoded by the encoder $\mathcal{E}_{2}$ and concatenated with the $z_{T}$ as the input of the U-Net, serves as the beginning of the reverse process. The employed 2D U-Net architecture adopts a multi-scale hierarchical structure comprising a time embedding module, an encoder module, a middle module, and a decoder module. This UNet model takes 8-channel input and produces 4-channel output, built upon a base channel configuration of 320. The encoder module begins with a 2D convolutional layer featuring a kernel size of 3 and padding of 1, followed by a series of residual blocks with progressively increasing feature dimensions. Each residual block includes multiple convolutional layers, GroupNorm [24] for normalization, the Swish activation function [25] for non-linearity, and optional dropout for regularization. Two residual blocks are employed per downsampling stage, with attention mechanisms activated at resolutions of 4, 2, and 1. The channel count increases across network layers following a multiplier sequence of 1, 2, 4, and 4. The decoder mirrors this structure, utilizing the same residual blocks to progressively upsample feature maps, reducing the number of channels in reverse order while merging them with corresponding encoder feature maps through skip connections. Additionally, the model incorporates 8 attention heads and employs a spatial transformer with a depth of 1, conditioned on a context dimension of 768.
Next, the target depth $d$ is embedded into the downsampling stage to allow the model to condition the latent representation with spatial hierarchy early in the process, which ensures that depth-related features, such as different tissue layers and bone structures, are incorporated from the outset, preserving the anatomical coherence of the volume. In the upsampling stage, the patient-specific radiographic feature information $r$ (refer to Table I) is embedded to guide the synthesis of fine clinical features based on the spatial hierarchy established during downsampling, such as the presence of cysts and the degree of attrition. Specifically, the depth $d$ is normalized to the range $[0,1]$. On the other hand, the patient-specific radiographic feature information $r$ is encoded using a one-hot encoding scheme.
TABLE I: Description of radiographic features
Variable Description
V00XROSFM Osteophytes (OARSI 0-3), femur medial compartment
V00XRSCFM Sclerosis (OARSI 0-3), femur medial compartment
V00XRCYFM Cysts (OARSI 0-1), femur medial compartment
V00XRJSM Joint space narrowing (OARSI 0-3), medial compartment
V00XRCHM Chondrocalcinosis (Grades 0-1), medial compartment
V00XROSTM Osteophytes (OARSI 0-3), tibia medial compartment
V00XRSCTM Sclerosis (OARSI 0-3), tibia medial compartment
V00XRCYTM Cysts (OARSI 0-1), tibia medial compartment
V00XRATTM Attrition (OARSI 0-1), tibia medial compartment
V00XRKL Kellgren and Lawrence (0-4) for osteoarthritis severity
V00XROSFL Osteophytes (OARSI 0-3), femur lateral compartment
V00XRSCFL Sclerosis (OARSI 0-3), femur lateral compartment
V00XRCYFL Cysts (OARSI 0-1), femur lateral compartment
V00XRJSL Joint space narrowing (OARSI 0-3), lateral compartment
V00XRCHL Chondrocalcinosis (Grades 0-1), lateral compartment
V00XROSTL Osteophytes (OARSI 0-3), tibia lateral compartment
V00XRSCTL Sclerosis (OARSI 0-3), tibia lateral compartment
V00XRCYTL Cysts (OARSI 0-1), tibia lateral compartment
V00XRATTL Attrition (OARSI 0-1), tibia lateral compartment
    OARSI: Osteoarthritis Research Society International Grading System
@@
$\mathop{\min}_{\theta}\mathbb{E}_{z\sim\varepsilon_{1}(x_{d}),t,\epsilon\sim\mathcal{N}(0,\mathbf{I})}||\epsilon-\epsilon_{\theta}[z_{t},t,(z_{T},y,d,r)]||_{2}^{2}$ (10)
Finally, for each reconstructed slice at the depth $d$, denoted as $\hat{x}_{d}$, it is generated through a decoder $\mathcal{D}$. During the inference phase, this process is repeated $S$ times, corresponding to the depth (i.e., channel) $S$ of the original MRI volume. The final generated MRI sequence $\hat{x}$ is formed by stacking all generated slices, represented as $\hat{x}=\{\hat{x}_{d}\}_{d=0}^{S-1}$. For more clarity, Algorithm 1 and Algorithm 2 describe the training and inference processes of the proposed approach, respectively.
Input: $x$, $t$, $y$, $r$, $S$, $\mathcal{E}_{2}$, and all initial parameters
Output: $\hat{x}_{d}$, $\mathcal{E}_{1}$, $\mathcal{D}$, $\mathcal{U}$ and learned parameters
@@
             Get $\epsilon_{\theta},\hat{z}_{d}\leftarrow\mathcal{U}(z_{T},\mathcal{E}_{2}(y),t,d,r)$
@@
Input: $t$, $r$, $y$, $S$, $\mathcal{E}_{2}$, $\mathcal{D}$ and $\mathcal{U}$
@@
       Get $\hat{z}_{d}\leftarrow\mathcal{U}(\epsilon,\mathcal{E}_{2}(y),t,d,r)$
@@
III Experimental settings
This section provides an overview of the experimental data, the preprocessing steps undertaken, and the specifics of the experimental procedures.
III-A Employed database